OpenStudy (anonymous):

Evaluate the surface integral. ∫∫S ((x^2)z + (y^2)z) dS S is the hemisphere x2 + y2 + z2 = 9, z ≥ 0.

OpenStudy (anonymous):

\[\int\limits_{}^{}\int\limits_{\S}^{}(x^2z+y^2z) dS\] S is the hemisphere \[x^2+y^2+z^2 = 9\] \[z \ge 0\]